The MYZ Company is considering the purchase of a machine that costs 100 thousand dollars, and which has a lifespan of only two years, after which it has a zero scrap value. This investment, if undertaken, will generate gross returns of 40 thousand dollars and 64 thousand dollars at the end of the first and second years, respectively, after deducting all the costs except depreciation and interest costs. Should the Company go ahead with this investment when the prevailing rate of interest is 6 percent? Explain your answer


1
Expert's answer
2021-05-30T14:18:44-0400

Solution:

We use NPV to analyze the profitability of an investment.

NPV = PV of future cashflows – PV of initial investment

Under NPV calculation, depreciation is irrelevant.

First discount each future cash flow in order to get the present value of each cash flow, and then sum those present values associated with each time period.


NPV ="\sum \frac{Ct}{(1+r)^{t} } - Co"

Where: Ct = Cashflow per period

             r = rate of interest

             t = time

          Co = Cash outflow (initial investment)

rate of interest = 6% or 0.06

Total future cashflows for the 2 years will be as follows:


= "\frac{40,000}{(1+0.06)^{1} } + \frac{64,000}{(1+0.06)^{2} }"


= 37,735.85 + 56,959.77 = 94,695.62


NPV = 94,695.62– 100,000 = -5,304.38

NPV = (5,304.38)


NPV is negative, which indicates that the projected earnings generated by the investment in the present are less than the expected costs, and hence will be unprofitable.

Therefore, MYZ company should not go ahead with this investment.